A point charge q = -0.25 nC is fixed at the origin.Where must a proton be placed in order for the electric force acting on it to be exactly opposite to its weight?(Let the y axis be vertical and the x axis be horizontal.)
r=? km
Answer #1
Charge q = -0.25 x 10 -9 CCharge of proton q ' = 1.6 x 10 -19 CMass of proton m = 1.67 x 10 -27 kgElectric force = weightKqq' / r 2 = mgFrom this required distance r = √[ Kqq' / mg ]Where K = Coulomb's constant= 8.99 x 10 9 Nm 2/ C2Substitue values we get r = 4687.47 m=4.68747 km
answered by: Assar